A clock has a pendulum 22 cm long. If it swings through an angle of 32 degrees, how far does the bottom of the bob travel in one swing?
Found 2 solutions by zarabella, Alan3354:
Answer by zarabella(1)

You need to find the length of the arc
L= arc ength
L=(q/180) * (pi) * (r) where q=32ยบ , pi=3.14159, and r=22 cm
L=(32/180)(pi)(22)
L=(8/45)(pi)(22)
L=(176/45)(pi) cm, or
L=(176/45)(3.14159)
L=552.91984/45
L=12.287 cm

Answer by Alan3354(69443)

A clock has a pendulum 22 cm long. If it swings through an angle of 32 degrees, how far does the bottom of the bob travel in one swing?
==============
Find the circumference of a circle with a radius of 22 cm.
Multiply that by 32/360
====================================
PS Pendulum clocks seldom, if ever, have a swing of 10 degrees or more off center.